Solving Word Problems Using Quadratic Equations
24. A rectangular field measures 15 m by 20 m. A rectangular area is to be fenced in by reducing each dimension by the same amount. The fenced-in area be?
Thanks@@!!!

Answer by nerdybill(7384)

24. A rectangular field measures 15 m by 20 m. A rectangular area is to be fenced in by reducing each dimension by the same amount. The fenced-in area be?
.
Let x = reduction
then
15-x = width
20-x = length
.
Area = (15-x)(20-x)
Area = (15-x)(20-x)
Area = 300-15x-20x-x^2
Area = 300-35x-x^2
Area = -x^2-35x+300